# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ 1 Introduction
Target tracking is a fundamental control task for autonomous agents, allowing them to be used in a variety of applications including environmental monitoring Aucone et al. (2023), agriculture Daponte et al. (2019), and air shows Schoellig et al. (2014) to name a few. Typically, an autonomous agent is given a reference trajectory, which should be tracked with as little error as possible. In the case of linear systems with quadratic costs the problem, also known as Linear Quadratic Tracking (LQT), can be formulated as follows. Let the autonomous agent be governed by the linear time-invariant dynamics
@@
where $x_{t}\in\mathbb{R}^{n}$ is the state, and $u_{t}\in\mathbb{R}^{m}$ is the control input, while $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are the known system dynamics. Given a target trajectory $r_{t}\in\mathbb{R}^{n\times 1}$, $t\geq 0$ the goal is finding the optimal control input of
@@
subject to the dynamics (1). Here, $T$ is the control horizon. The state, input, and terminal penalties, $Q\in\mathbb{R}^{n\times n},R\in\mathbb{R}^{m\times m},Q_{T}\in\mathbb{R}^{n\times n}$, respectively, are design choices.
In classical target tracking, the target state $r_{t}$ is known a priori and is precomputed. However, in many cases, the target might be time-varying and unknown. Such scenarios arise, for example, in the case of adversarial target tracking, wild animal tracking, moving obstacle avoidance, pedestrian tracking, etc. In such cases, the target trajectory is generated by an external unknown dynamical system, also known as “exosystem” Nikiforov and Gerasimov (2022). In this work, we assume that the target is generated by time-varying, auto-regressive (AR) dynamics. Let
@@
Adaptive control has a long history of dealing with dynamical uncertainty Annaswamy and Fradkov (2021), where the goal is to simultaneously control the system and adapt to the uncertainty. One of the most widely-used algorithms in the adaptive control literature with ubiquitous applications has been the celebrated Recursive Least Squares (RLS) algorithm with forgetting factors Åström et al. (1977). While adaptive control has been extensively studied in the stochastic or time-invariant regime Guo and Ljung (1995); Ljung and Gunnarsson (1990), results for non-stochastic, time-varying systems are relatively scarce. By studying the problem through the modern lens of online learning and non-stochastic control Hazan and Singh (2022), we can provide guarantees for new classes of time-varying systems.
Most works on adaptive control follow one of the two possible paradigms: i) direct control and ii) indirect control, mirroring model-free and model-based Reinforcement Learning. In the former case, adaptation occurs directly at the control policy; in the latter case, a model of the uncertainty is kept at all times and the control policy is updated indirectly based on the current model. A notable benefit of indirect architectures is that they decouple control from learning, making it easier to incorporate changes in the control objective or robustness specifications. In the case of non-stochastic control, learning the optimal feedforward law Agarwal et al. (2019a); Foster and Simchowitz (2020) fits the direct paradigm. Prior works that use predictions Li et al. (2019); Zhang et al. (2021b) are closer to the indirect paradigm but they either do not provide a prediction method or the regret guarantees are suboptimal in the case of targets with dynamic structure.
@@ 1.1 Contribution
In this work, we propose a tracking control algorithm, called predictive linear online tracking (PLOT). It adapts to unknown targets online using sequentially measured target data, based on the indirect paradigm. PLOT uses a modified version of the RLS algorithm with forgetting factors Yuan and Lamperski (2020) to learn the target dynamics and predict future target states. Then, it computes a receding horizon control input using the predicted target states in a certainty equivalent fashion. To characterize closed-loop performance, we use the notion of dynamic regret Zinkevich (2003). In our setting, dynamic regret compares the incurred closed-loop cost of any online algorithm against the cost incurred by the optimal control actions in hindsight, that is, the actions generated by the optimal non-causal policy that has full knowledge about future target states. Our contributions are the following.
@@
Dynamic regret for online tracking. We prove that the dynamic regret of PLOT is upper-bounded by $\mathcal{O}(\sqrt{TV_{T}})$, where $V_{T}$ is the total variation of the target dynamics $S_{t}$ (path length). When the target dynamics are static (zero variation), the regret becomes logarithmic and we recover prior results Foster and Simchowitz (2020). While prior dynamic regret bounds exist for linear quadratic control, they either focus on direct online control Baby and Wang (2022), where we learn over a directly parameterized disturbance affine feedback policy, or they assume no structure for the targets Li et al. (2019); Karapetyan et al. (2023). Instead, we focus on indirect online control, where we learn an internal model representation for the unknown target. By setting up multiple step ahead predictors, we have more flexible and frequent feedback, thus avoiding extra logarithmic factors due to delayed learning feedback as in Baby and Wang (2022); Foster and Simchowitz (2020). Using the total variation of the target dynamics $S_{t}$ to characterize dynamic regret is a new point of view, improving prior bounds that use the variation of the target state $r_{t}$ itself Li et al. (2019); Karapetyan et al. (2023).
@@
Prediction of time-varying partially observed systems. We employ the RLS algorithm to predict target states multiple time steps into the future. Our result is of independent interest since it applies to time-varying AR systems or systems with exogenous inputs (ARX), which is a common model in system identification. We obtain dynamic regret guarantees for prediction by adapting the analysis of Yuan and Lamperski (2020) to the setting of learning with delayed feedback Joulani et al. (2013). This result can be seen as a generalization of time-invariant Kalman filtering to the non-stochastic, time-varying, multi-step ahead case.
@@
Experimental demonstration. While the theory of online non-stochastic control has matured over the past years Hazan and Singh (2022), its practice is still lagging. Despite promising results Suo et al. (2021), most demonstrations are limited to simulation environments. In this work, we perform extensive comparisons of several online control methods in simulations and implement PLOT on a hardware setup. To facilitate benchmarking and future developments, we make our code open-source and implement it on Crazyflie drones with a reconfigurable software architecture built on the Robot Operating System (ROS). To the best of our knowledge, our work is one of the first to demonstrate the use of online non-stochastic control with guarantees on a real robot.
1.2 Related Work
Online and non-stochastic control.
Online and non-stochastic control techniques have been applied to systems with known Agarwal et al. (2019a) as well as unknown system dynamics Simchowitz (2020). In this paper, we focus on the case of known system dynamics, where the main purpose of the online controller is to react to (non-stochastic) unknown disturbances/targets. [^2] Note that as shown in many works, see Foster and Simchowitz (2020); Karapetyan et al. (2023), the problem of tracking is equivalent under certain conditions to the problem of disturbance rejection. Hence, for the purpose of this literature review, we can use disturbance/target interchangeably. In this setting, Agarwal et al. (2019a) achieved $\mathcal{O}(\sqrt{T})$ regret for general convex stage cost functions under a disturbance affine feedback policy. In the case of quadratic costs, and more generally strongly convex stage costs [^3] Certain additional conditions are required., the guarantees were improved to logarithmic Foster and Simchowitz (2020); Simchowitz (2020); Agarwal et al. (2019b). The aforementioned works consider static problems in the sense that they compete against the best stationary policy. In Baby and Wang (2022) dynamic regret guarantees for the quadratic cost case (LQR) were derived using the online Newton step algorithm along with the Follow the Leading History (FLH) algorithm. It was proved that the dynamic regret with respect to any dynamic disturbance affine policy is upper bounded by $\mathcal{O}(T^{1/3}V^{2/3}_{T})$, where $V_{T}$ is the total variation of the parameters of the policy. Dynamic regret guarantees of the order of $\mathcal{O}(\sqrt{TV_{T}})$ were also derived in Zhao et al. (2022) for the same setting, using gradient-based (first order) base learners. Learning disturbance affine feedback policies follows the direct online control paradigm since we adapt the policy parameters directly.
In this paper, we study implicit online control architectures, where we decouple online prediction of disturbances/targets from control design. Several works have studied the effect of predictions on receding horizon control Yu et al. (2020); Li et al. (2019); Zhang et al. (2021b). However, either no method of prediction is provided or the dynamic regret guarantees are suboptimal in our setting. In particular, the regret guarantees scale with the total variation of the disturbance/target itself $\sum_{t=0}^{T-1}\|r_{t+1}-r_{t}\|$ Li et al. (2019), which is appropriate for unstructured or slowly-varying targets but is suboptimal in the case of targets with dynamic structure; there exist targets for which the total variation of their state $r_{t}$ is linear, while the total variation of the target dynamics $S_{t}$ is zero–see Appendix B.1.
Abbasi-Yadkori et al. (2014) introduced the problem of non-stochastic control for target tracking, where the target can be adversarial. Logarithmic regret guarantees are provided for learning affine control policies with constant affine terms. However, policies with constant affine terms can only track static points as targets; they are not rich enough to track trajectories that are not points and, in general, time-varying targets. Recently, Niknejad et al. (2023) studied online learning for target tracking in the case of unknown system dynamics and unknown time-invariant target dynamics. Our setting is different since we have known dynamics for the system, time-varying dynamics for the target, and we use dynamic regret. Their static regret guarantee of $T^{2/3}$ is conservative for our setting, where we can obtain logarithmic regret in the static case, however, it applies to general convex costs and unknown dynamics. In Karapetyan et al. (2023), dynamic regret guarantees are provided for target tracking, under a first order method which treats the targets as slowly varying or steady-state. Similar to Li et al. (2019), the dynamic regret depends on the total variation of the target state which can be suboptimal in the case of target with dynamics. Zhang et al. (2022) proved $\tilde{\mathcal{O}}(\sqrt{|\mathcal{I}|})$ (for any interval $\mathcal{I}$) adaptive regret guarantees for tracking in the case of convex costs. Their setting is different since no dynamic structure is assumed for the targets, the comparator class is controllers which are constant over the interval of interest, while we focus on square losses and dynamic regret.
Although related, our setting is different from Gradu et al. (2023) and Minasyan et al. (2021) where the system itself is linear time varying, there is no target, and adaptive regret is considered as a performance metric. Instead, here, the system’s dynamics are time invariant, the target’s dynamics are unknown and time varying, and we consider dynamic regret. Safe online non-stochastic control subject to constraints has also been studied before Zhou et al. (2023); Nonhoff et al. (2023); Li et al. (2021). Dealing with constraints is a more challenging problem in general. In many cases, the analysis is restricted to the static case Li et al. (2021), requires relaxing the notion of policy regret Zhou et al. (2023) or the path length scales with the target state Nonhoff et al. (2023), which is not always optimal in our setting as we discuss above.
Finally, we note that when the system dynamics are unknown, it is no longer possible to achieve logarithmic regret in the general case, even under stochastic disturbances Simchowitz and Foster (2020); Ziemann and Sandberg (2022). It is possible to achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret Chen and Hazan (2021); Simchowitz (2020); Mania et al. (2019) instead.
Prediction and online least squares.
Online linear regression and the least squares method in particular have been studied extensively before Azoury and Warmuth (2001); Cesa-Bianchi and Lugosi (2006). Anava et al. (2013) provided logarithmic regret guarantees for online prediction of auto-regressive (AR) and auto-regressive with moving average (ARMA) systems using the online Newton step (ONS) method. The result holds for exp-concave prediction loss functions, which includes the square prediction loss. Similar guarantees were proved for partially observed stochastic linear systems in the case of Kalman filtering Ghai et al. (2020); Tsiamis and Pappas (2022); Rashidinejad et al. (2020). Under first-order methods, the regret guarantees become $\mathcal{O}(\sqrt{T})$ Anava et al. (2013); Kozdoba et al. (2019), however, first-order methods are also known to adapt to changes Zinkevich (2003); Hazan and Seshadhri (2009). On the contrary, the standard versions of second-order methods like RLS or ONS are not adequate for time-varying settings, where adaptation is needed. Another challenge is that in the case of multi-step ahead prediction, these methods should be adapted to deal with delayed learning feedback.
To solve the former issue, one way is to employ multiple learners along with a meta-algorithm that chooses the best expert online, also known as the Follow the Leading History (FLH) algorithm Hazan and Seshadhri (2009); Baby and Wang (2022). Another option is to employ exponential forgetting Yuan and Lamperski (2020); Ding et al. (2021). In this paper, we follow the second approach since it has been standard practice in the adaptive control literature Guo and Ljung (1995); Åström et al. (1977). Moreover, it has a simple implementation for real-time applications, like drone control.
Control Literature.
Common feedback control policies, such as the proportional-integral-derivative (PID) controllers find extensive use for tracking problems in practice Pan et al. (2018); Cervantes and Alvarez-Ramirez (2001). For example, integral control has been used to track constant unknown offsets. However, they typically require fine-tuning and their tracking performance for unknown time-varying trajectories is not guaranteed in general. Tracking of known reference trajectories with uncertain repetitive dynamics is studied in the context of iterative learning control Bristow et al. (2006); Owens and Hätönen (2005) with online extensions studying dynamic regret Balta et al. (2022). However, the methods do not extend to non-repetitive setting of LQT with unknown time-varying references.
Predictive controllers such as model predictive control (MPC) have also been used to track constant references, see, for example, offset-free MPC Pannocchia and Bemporad (2007); Maeder et al. (2009). Various tracking MPC methods are proposed for time-varying references, e.g., Köhler et al. (2020). However, unless we have an accurate prediction model for the reference trajectory, the guarantees do not extend to more general sequentially revealed unknown trajectories. Dual MPC methods with active exploration are proposed for tracking applications and uncertain dynamics Heirung et al. (2017); Soloperto et al. (2019); Parsi et al. (2022). However, transient performance, regret guarantees, or the case of unknown time-varying references are not studied.
The LQT problem with unknown targets has been studied before from the point of view of control theory and adaptive control Vamvoudakis (2016); Modares and Lewis (2014); Peterson and Narendra (1982). Typically, the goal is to prove asymptotic convergence or boundedness of tracking error by appealing to Lyapunov stability theory. Most results assume time-invariant target dynamics. Tracking time-varying targets is possible using adaptive control techniques, under the assumption that the parameters of the target evolve like a random walk Guo and Ljung (1995). However, this assumption excludes adversarial targets. By using regret as a metric, we can obtain non-asymptotic guarantees that also capture adversarial, non-stochastic behaviors.
Finally, an alternative approach is to employ robust control techniques to account for worst-case disturbances/tracking errors. Notable approaches include the celebrated $\mathcal{H}_{\infty}$ control approach or mixed $\mathcal{H}_{2}-\mathcal{H}_{\infty}$ control Zhang et al. (2021a). Recently, robust controllers inspired by the notion of regret were designed Goel and Hassibi (2023).
@@ 2 Problem Statement
Many fundamental targets can be captured by dynamics (3). For example, a target moving on a circle or a straight line with a constant speed can be captured by linear time-invariant dynamics–see Appendix B.1. By considering time-varying dynamics, we can model more complex target trajectories, e.g., combinations of lines, circles, switching patterns, etc. We remark that the LQT problem is equivalent to the Linear Quadratic Regulator (LQR) with disturbances–see Appendix A. Hence, the techniques applied here could also be applied to noisy system dynamics (1), where we attach a dynamic structure to the disturbances. Our results can be generalized directly to targets driven by exogenous variables (ARX) ${r_{t+1}=S_{t+1}z_{t}+v_{t+1}}$–it is sufficient to redefine $\tilde{z}_{t}=\begin{bmatrix}z^{\top}_{t}&1\end{bmatrix}^{\top}$ and $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$. In this case, we can capture richer targets, including other control systems.
@@
where $k\in\mathbb{N}$ is the number of future steps. The $k-$step ahead matrix $S_{t+k|t}$ is a nonlinear function (multinomial) of $S_{t+k},\dots,S_{t+1}$. Expressions for the multi-step ahead recursions (4) can be found in Appendix B.2. By definition $S_{t+1|t}\equiv S_{t+1}$. The notation $S_{t+k|t}$ indicates that given all information $z_{t}$ at time $t$, the product $S_{t+k|t}z_{t}$ acts as a $k-$step ahead prediction of $r_{t+k}$ at time $t$.
@@
Such boundedness conditions reflect typical assumptions in online learning Yuan and Lamperski (2020). The second condition allows general $S_{k}$ as long as the multi-step ahead matrices do not blow-up.
@@ 2.1 Control Objective
Our goal is to design an online controller that adapts to the unknown target dynamics. To evaluate online performance, we use dynamic regret, which compares the incurred cost of the online controller to the optimal controller in hindsight that knows all future target states in advance (called simply the optimal controller). For a target state realization $\textbf{r}:=r_{0:T}$, we denote the cumulative cost by
$J_{T}(u_{0:T-1};\textbf{r})\triangleq\sum_{t=0}^{T-1}(\lVert x_{t}-r_{t}\rVert^{2}_{Q}+\lVert u_{t}\rVert^{2}_{R})+\lVert x_{T}-r_{T}\rVert_{Q_{T}}^{2}.$
@@
$\mathcal{R}(\pi)=J_{T}(u_{1:T-1}^{\pi};\textbf{r})-J_{T}(u_{1:T-1}^{*};\textbf{r}).$ (5)
We can now summarize the main objective of the paper.
@@
To make sure that the control problem is well-defined, we introduce the following assumption which is standard in the control literature Zhang et al. (2021b). It guarantees that the LQT controller will be (internally) stable.
@@
We assume the terminal cost matrix $Q_{T}$ is chosen as the solution $P$ of the Discrete Algebraic Riccati Equation.
@@
We select $Q_{T}=P$, where $P$ is the unique solution to Riccati equation
$P=Q+A^{\top}PA-A^{\top}PB(R+B^{\top}PB)^{-1}B^{\top}PA.$ (6)

### v3

@@ Abstract
Online Learning, Tracking, Control, Recursive Learning
[…]
skipabove=6pt, skipbelow=0pt
@@ 1 Introduction
Target tracking is a fundamental control task for autonomous agents, allowing them to be used in a variety of applications including environmental monitoring (Aucone et al., 2023), agriculture (Daponte et al., 2019), and air shows (Schoellig et al., 2014) to name a few. Typically, an autonomous agent is given a reference trajectory, which should be tracked with as little error as possible. In the case of linear systems with quadratic costs the problem, also known as Linear Quadratic Tracking (LQT), can be formulated as follows. Let the autonomous agent be governed by the linear dynamics
@@
where $x_{t}\in\mathbb{R}^{n}$ is the state, and $u_{t}\in\mathbb{R}^{m}$ is the control input, while $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ denote the known system dynamics. Given a target trajectory $r_{t}\in\mathbb{R}^{n\times 1}$, $t\geq 0$ the goal is finding the optimal control input of the finite-horizon problem
@@
subject to the dynamics (1), for some total time (horizon) $T$. The state, input, and terminal penalties, $Q\in\mathbb{R}^{n\times n},R\in\mathbb{R}^{m\times m},Q_{T}\in\mathbb{R}^{n\times n}$, respectively, are design choices.
In classical target tracking, the target state $r_{t}$ is known a priori and is precomputed. However, in many cases, the target might be time-varying and unknown. Such scenarios arise, for example, in the case of adversarial target tracking, wild animal tracking, moving obstacle avoidance, pedestrian tracking, etc. In such cases, the target trajectory is generated by an external unknown dynamical system, also known as “exosystem” (Nikiforov & Gerasimov, 2022). In this work, we assume that the target is generated by time-varying, auto-regressive (AR) dynamics. Let
@@
Adaptive control has a long history of dealing with dynamical uncertainty (Annaswamy & Fradkov, 2021), where the goal is to simultaneously control the system and adapt to the uncertainty. One of the most widely-used algorithms in the adaptive control literature with ubiquitous applications has been the celebrated Recursive Least Squares (RLS) algorithm with forgetting factors (Åström et al., 1977). While adaptive control has been extensively studied in the stochastic or time-invariant regime (Guo & Ljung, 1995; Ljung & Gunnarsson, 1990), results for non-stochastic, time-varying systems are relatively scarce. By studying the problem through the modern lens of online learning and non-stochastic control (Hazan & Singh, 2022), we can provide guarantees for new classes of time-varying systems.
Most works on adaptive control follow one of the two possible paradigms: i) direct control and ii) indirect control, mirroring model-free and model-based Reinforcement Learning. In the former case, adaptation occurs directly at the control policy; in the latter case, a model of the uncertainty is kept at all times and the control policy is updated indirectly based on the current model. A notable benefit of indirect architectures is that they decouple control from learning, making it easier to incorporate changes in the control objective or robustness specifications. In the case of non-stochastic control, learning the optimal feedforward law (Agarwal et al., 2019a; Foster & Simchowitz, 2020) fits the direct paradigm. Prior works that use predictions (Li et al., 2019; Zhang et al., 2021b) are closer to the indirect paradigm but they either do not provide a prediction method or the regret guarantees are suboptimal in the case of targets with dynamic structure.
@@ 1.1 Contribution
In this work, we propose a tracking control algorithm, called Predictive Linear Online Tracking (PLOT). It adapts to unknown targets online using sequentially measured target data, based on the indirect paradigm. PLOT uses a modified version of the RLS algorithm with forgetting factors (Yuan & Lamperski, 2020) to learn the target dynamics and predict future target states. Then, it computes a receding horizon control input using the predicted target states in a certainty equivalent fashion. To characterize closed-loop performance, we use the notion of dynamic regret (Zinkevich, 2003). In our setting, dynamic regret compares the incurred closed-loop cost of any online algorithm against the cost incurred by the optimal control actions in hindsight, that is, the actions generated by the optimal non-causal policy that has full knowledge about future target states. Our contributions are the following.
Dynamic regret for online tracking.
@@
We prove that the dynamic regret of PLOT is upper-bounded by $\mathcal{O}(\sqrt{TV_{T}})$, where $V_{T}$ is the total variation of the target dynamics $S_{t}$ (path length). When the target dynamics are static (zero variation), the regret becomes logarithmic and we recover prior results (Foster & Simchowitz, 2020). While prior dynamic regret bounds exist for linear quadratic control, they either focus on direct online control (Baby & Wang, 2022), where we learn over a directly parameterized disturbance affine feedback policy, or they assume no structure for the targets (Li et al., 2019; Karapetyan et al., 2023). Instead, we focus on indirect online control, where we learn an internal model representation for the unknown target. By setting up multiple step ahead predictors, we have more flexible and frequent feedback, thus avoiding extra logarithmic factors due to delayed learning feedback as in (Baby & Wang, 2022; Foster & Simchowitz, 2020). Using the total variation of the target dynamics $S_{t}$ to characterize dynamic regret is a new point of view, improving prior bounds that use the variation of the target state $r_{t}$ itself (Li et al., 2019).
Prediction of time-varying partially observed systems.
@@
We employ the RLS algorithm to predict target states multiple time steps into the future. Our result is of independent interest since it applies to time-varying AR systems or systems with exogenous inputs (ARX), which is a common model in system identification. We obtain dynamic regret guarantees for prediction by adapting the analysis of Yuan & Lamperski (2020) to the setting of learning with delayed feedback (Joulani et al., 2013). This result can be seen as a generalization of time-invariant Kalman filtering to the non-stochastic, time-varying, multi-step ahead case.
Experimental demonstration.
@@
While the theory of online non-stochastic control has matured over the past years (Hazan & Singh, 2022), its practice is still lagging, with only a few notable experimental demonstrations on hardware (Suo et al., 2021; Snyder et al., 2023). In this work, we perform extensive simulations showing how the regret analysis can be used as a tool to tune the controller hyperparameters, provide comparisons of several online control methods in simulations, and implement PLOT on a hardware setup. To facilitate benchmarking and future developments, we make our code open-source and implement it on Crazyflie drones with a reconfigurable software architecture built on the Robot Operating System (ROS). To the best of our knowledge, our work is one of the first to demonstrate the use of online non-stochastic control with guarantees on a real quadrotor.
@@ 2 Problem Statement
Many fundamental targets can be captured by dynamics (3). For example, a target moving on a circle or a straight line with a constant speed can be captured by linear time-invariant dynamics–see Appendix C.1. By considering time-varying dynamics, we can model more complex target trajectories, e.g., combinations of lines, circles, switching patterns, etc. Our results can be generalized directly to targets driven by exogenous variables (ARX) ${r_{t+1}=S_{t+1}z_{t}+v_{t+1}}$–it is sufficient to redefine $\tilde{z}_{t}=\begin{bmatrix}z^{\top}_{t}&1\end{bmatrix}^{\top}$ and $\tilde{S}_{t}\triangleq\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$. We can thus capture even richer targets, including other control systems. In fact, any arbitrary bounded adversarial target can be captured by such a model; it is sufficient to take $S_{t+1}=0$ and $v_{t+1}=r_{t+1}$. We note that the above representation is not unique–see also Appendix C.1.
We remark that the LQT problem is equivalent to the Linear Quadratic Regulator (LQR) with disturbances–see Appendix B. Hence, the techniques applied here could also be applied to noisy system dynamics (1), where we attach a dynamic structure to the disturbances.
@@
where $k\in\mathbb{N}$ is the number of future steps. The $k-$step ahead matrix $S_{t+k|t}$ is a nonlinear function (multinomial) of $S_{t+k},\dots,S_{t+1}$. Expressions for the multi-step ahead recursions (4) can be found in Appendix C.2. By definition $S_{t+1|t}\equiv S_{t+1}$. The notation $S_{t+k|t}$ indicates that given all information $z_{t}$ at time $t$, the product $S_{t+k|t}z_{t}$ acts as a $k-$step ahead prediction of $r_{t+k}$ at time $t$.
@@
Such boundedness conditions reflect typical assumptions in online learning (Yuan & Lamperski, 2020). The second condition allows general $S_{k}$ as long as the multi-step ahead matrices do not blow up.
@@ 2.1 Control Objective
Our goal is to design an online controller that adapts to the unknown target dynamics. To evaluate the online performance, we use dynamic regret, which compares the incurred cost of the online controller to the optimal controller in hindsight that knows all future target states in advance (called simply the optimal controller). For a target state realization $\textbf{r}:=r_{0:T}$, we denote the cumulative cost by
$J_{T}(u_{0:T-1};\textbf{r})\!\triangleq\!\sum_{t=0}^{T-1}\!\left(\lVert x_{t}\!-\!r_{t}\rVert^{2}_{Q}\!+\!\lVert u_{t}\rVert^{2}_{R}\right)+\lVert x_{T}-r_{T}\rVert_{Q_{T}}^{2}.$
@@
$\mathcal{R}(\pi)=J_{T}(u_{1:T-1}^{\pi};\textbf{r})-J_{T}(u_{1:T-1}^{*};\textbf{r}),$ (5)
defined for a target state realization r. We can now summarize the main objective of the paper.
{mdframed}
[roundcorner=3pt, backgroundcolor=blue!10]
@@
To make sure that the control problem is well-defined, we introduce the following assumption which is standard in the control literature (Zhang et al., 2021b). It guarantees that the LQT controller will be (internally) stable.
@@
We assume the terminal cost matrix $Q_{T}$ is chosen as the solution $X$ of the Discrete Algebraic Riccati Equation.
@@
We select $Q_{T}=X$, where $X$ is the unique solution to Riccati equation
$X=Q+A^{\top}XA-A^{\top}XB(R+B^{\top}XB)^{-1}B^{\top}XA.$ (6)